The circuit to the right consists of a battery (Vo-31.5 V) and five resistors (R1 711 Ω R2-382 Ω , R3 163 Ω , R4-634 Ω , and R5-565 Ω). Find the current passing through each of the specified points. Number mA B C R. Number MA Number MA
Answer #1

net resistance of the circuit

R = 711 + ( 382^-1 + 163^-1 + 634^-1 + 565^-1)^-1

R = 793.644 ohm

net current through circuit

i = V/R = 31.5 / 793.644 = 39.69 mA

voltage across R2 = R3 = R4 = R5

V' = 31.5 - 0.03969* 711

V' = 3.28 V

=========

current through D

id = V' / R5 = 3.28 / 565 = 5.805 mA

=======

current through A = 3.28/ 382 = 8.586 mA

current through R3 = 3.28/ 163 = 20 mA

Current through B = 8.586 + 20 = 28.586 mA

=======

current through F = i - current through A

Current through F = 39.69 - 8.586 = 31.1 mA

======

do comment in case any doubt, will reply for sure. Goodluck